12. Determine the derivative of the function. f(x) = (x - 1)e7x + 6 a. f(x) = 27x+6(7x – 6) b. $(x) = 27x+617x + 6) C. $(x
Answer #1

$1x) = (x-1) (1948) Applying froduct rule of differentation dbal= (x-1) 1 Ento) + (+6)xi f als eToto [ 74-7 ti] [8/4)= 4T+6